# Channel-State duality with centers

### v1

@@ I Introduction
The focus in this work is on mappings encoding information transport properties in quantum systems, specifically those which have an algebra of observables possessing nontrivial center, and, a representation of the same algebra on Hilbert spaces with a direct sum structure[16, 13, 17]. These mappings can also be seen as way to characterize entanglement between subsets of quantum degrees of freedom of the same systems, in a context in which no unambiguous notion of subsystem is immediately available, due to the Hilbert space lacking a tensor product decomposition.
Let us give a simple example of the maps we consider, but in a much simpler context, i.e. the simplest bipartite quantum system made from 2 qubits. For Bell states of 2 qubits, the action of an operator on one of the qubits can be expressed through an operator on the other, more specifically its transpose: [^1] For the |$\ket{\Psi^{\pm}}$ states, the corresponding operator is $X\sigma^{1}$, so multiplied by a Pauli matrix.
@@
which takes any (bounded) operator on a subsystem and turns it into an operator on its complementary subsystem. Such a mapping is state-dependent. How well subsystems are mapped onto each other by its action depends both on how similar the subsystems are (in particular, the dimensions of the subsystem or their tensor product substructure) and how strongly correlated they are in the given state.
Given any operator $X\in\mathbb{B}(\mathbb{C}^{2}_{A})$, we can write it, using a notation that applies straightforwardly in the general case, as
@@
One can verify easily in components that this does indeed give the transpose ±$\pm X^{t}$ of the operator $X$ for |$\ket{\Phi^{\pm}}$. Written like this, it becomes clear that this is a particular case of Channel-state duality[13, 17].
The duality connects sets of $k$-positive operators on a bipartite system $\mathbb{H}=\mathbb{H}_{A}\otimes\mathbb{H}_{B}$,
@@
$\begin{gathered}BL_{k}(\mathbb{B}(\mathbb{H}_{A}),\mathbb{B}(\mathbb{H}_{B}))=\{T:\mathbb{B}(H_{A})\rightarrow\mathbb{B}(H_{B}):\\ \mathbb{I}_{\mathbb{C}^{k}}\otimes T\text{ positive}\}\end{gathered}$ (5)
@@
In particular, it is a way to characterize quantum correlations beyond the case where the subsystems have identical Hilbert spaces, or when the total Hilbert space does not factorise at all in terms of subsystem Hilbert spaces.
Superoperators of this kind have found use in tensor network models of holography[5, 11, 6, 29], where the systems $A$ and $B$ are seen as, respectively, the bulk and boundary of a spatially compact system, and state or operator reconstruction on boundaries are used as hallmarks of holographic behaviour. Indeed, this is a case in which the relevant Hilbert spaces are not, in general, composed of factors of the same size when they factorize over $A$ and $B$, or do not factorize at all [7, 10, 3]. Motivated by these models, we study transport superoperators on bi- and tripartite systems subject to constraints which prevent the total Hilbert space to fully factorize, enforcing a structure of the form
$\mathbb{H}=\bigoplus_{E}\mathbb{H}_{I,E}\otimes\mathbb{H}_{O,E}.$ (10)
@@
The structure of this paper is as follows. In section II, we review the problem and introduce relevant concepts (e.g. algebraic subsystems, extension and partial trace maps). In section III, we study the case of transport operators on tripartite systems with no direct sum, and then extend the discussion to direct sums of tripartite systems. Section IV presents the generalization of the Channel-State duality to the direct sum case. Section V discusses possible extensions of these notions to the infinite dimensional case.
II The setting
@@
III Mappings on algebras with nontrivial centers
@@ III.2 Nontrivial center
IV Channel-State duality
@@ Acknowledgements
SL would like to thank D. Aliverti and L.Csillag for valuable feedback on the manuscript and useful discussions. SL and DO acknowledge funding from the Munich Center for Quantum Science and Technology. DO also acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG).

### v2

@@ I Introduction
In quantum information theory, the channel-state duality [1, 2] refers to a correspondence between quantum operations (specifically, quantum channels) and quantum states of composite systems, which has profound implications at both mathematical and physical levels. To be more specific, consider two systems, $A$ and $B$, described by Hilbert spaces $\mathbb{H}_{A}$ and $\mathbb{H}_{B}$, respectively, and a quantum channel $\mathcal{C}$ that maps states $\rho_{A}\in\mathbb{B}(\mathbb{H}_{A})$ to states $\rho_{B}\in\mathbb{B}(\mathbb{H}_{B})$. The channel-state duality establishes a correspondence between $\mathcal{C}$ and a specific state $\rho_{\mathcal{C}}\in\mathbb{B}(\mathbb{H})$ over the extended Hilbert space $\mathbb{H}=\mathbb{H}_{A}\otimes\mathbb{H}_{B}$. The state $\rho_{\mathcal{C}}$ is constructed by applying $\mathcal{C}$ to a branch of a maximally entangled state $\rho\in\mathbb{B}(\mathbb{H}_{A}\otimes\mathbb{H}_{A})$, as depicted in figure 1.
In this setup, if $\rho_{A}$ represents an initial state and $\rho_{B}$ the final state, the quantum channel can be viewed as implementing a form of “discrete dynamics”. In this perspective, the channel-state duality relates the kinematic properties of states to the dynamical properties of channels. Another relevant perspective, which is central to our work, arises from the fact that quantum states inherently define a quantum channel between two of their subsystems, as emphasized e.g. in [3].
Figure 1: A quantum channel $\mathcal{C}$, which maps states from system $A$ to system $B$, is dual to a state $\rho_{\mathcal{C}}$ of the composite system $A+B$. This state is constructed by applying $\mathcal{C}$ to one half of a maximally entangled state $\rho$ shared between two copies of system A.
In this work, we explore the channel-state duality in cases where the Hilbert space of the composite system $\mathbb{H}$ does not factorize over subsystems, but features a direct sum structure: $\mathbb{H}=\bigoplus_{\mu}\mathbb{H}^{\mu}_{A}\otimes\mathbb{H}^{\mu}_{B}$. This corresponds to situations where the algebras of observables for the systems possess a non-trivial center. In such cases, the channel-state duality can be used to characterize entanglement between subsets of quantum degrees of freedom, despite the fact that the Hilbert space lacks a clear tensor product decomposition, and so no unambiguous notion of subsystems is immediately available. Such Hilbert space structures generically arise in systems which are subject to holonomic constraints, but are also of importance in other contexts.
In particular, a Channel-state duality of this more general kind has found use in tensor network models of holography[4, 5, 6, 7], where the systems $A$ and $B$ are seen as, respectively, the bulk and boundary of a spatially compact system, and state or operator reconstruction on boundaries are used as hallmarks of holographic behaviour. Indeed, this is a case in which the relevant Hilbert spaces are not, in general, composed of factors of the same size when they factorize over $A$ and $B$, or do not factorize at all [8, 9, 10]. Motivated by these models, we study transport superoperators on bi- and tripartite systems subject to constraints which prevent the total Hilbert space to fully factorize, enforcing a structure of the form
$\mathbb{H}=\bigoplus_{E}\mathbb{H}_{I,E}\otimes\mathbb{H}_{O,E}.$ (1)
We will provide here a general framework to address these questions, leaving specific applications to future work.
The structure of this paper is as follows. In section II, we recapitulate the exact statement of Channel-state duality and present examples of non-factorised Hilbert spaces. In section III, we review the problem and introduce relevant concepts (e.g. algebraic subsystems, extension and partial trace maps). In section IV, we study the case of transport operators on tripartite systems with no direct sum, and then extend the discussion to direct sums of tripartite systems. Section V presents the generalization of the Channel-State duality to the direct sum case. Section VI discusses possible extensions of these notions to the infinite dimensional case.
II Channel-state duality
Let us first give a simple example of the maps we consider, but in a much simpler context, i.e. the simplest bipartite quantum system made from 2 qubits. For Bell states of 2 qubits, the action of an operator on one of the qubits can be expressed through an operator on the other, more specifically its transpose: [^1] For the |$\ket{\Psi^{\pm}}$ states, the corresponding operator is $X\sigma^{1}$, so multiplied by a Pauli matrix.
@@ II Channel-state duality
which takes any (bounded) operator on a subsystem and turns it into an operator on its complementary subsystem. Specifically,
@@
One can verify easily in components that this does indeed give the transpose ±$\pm X^{t}$ of the operator $X$ for |$\ket{\Phi^{\pm}}$. Written like this, it becomes clear that this is a particular case of channel-state duality [1, 2]. Note that such a mapping is state-dependent. How well subsystems are mapped onto each other by its action depends both on how similar the subsystems are (in particular, the dimensions of the subsystem or their tensor product substructure) and how strongly correlated they are in the given state.
To be even more precise, the duality connects sets of $k$-positive operators on a bipartite system $\mathbb{H}=\mathbb{H}_{A}\otimes\mathbb{H}_{B}$,
@@
$\begin{gathered}BL_{k}(\mathbb{B}(\mathbb{H}_{A}),\mathbb{B}(\mathbb{H}_{B}))=\{\mathcal{T}:\mathbb{B}(H_{A})\rightarrow\mathbb{B}(H_{B}):\\ \mathbb{I}_{\mathbb{C}^{k}}\otimes\mathcal{T}\text{ positive}\}\end{gathered}$ (6)
@@
With this in mind, we want to now present some examples of Hilbert spaces with a direct sum structure resulting from a nontrivial center, before discussing what impedes usual channel-state duality.
@@
III The setting
@@
IV Mappings on algebras with nontrivial centers
@@ IV.2 Nontrivial center
V Channel-State duality
@@ Acknowledgements
SL would like to thank D. Aliverti and L.Csillag for valuable feedback on the manuscript and useful discussions. SL and DO acknowledge funding from the Munich Center for Quantum Science and Technology. DO also acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG), and the ATRAE programme of the Spanish Government, through the grant PR28/23 ATR2023-145735. ECs participation in this project was made possible by a DeBenectis Postdoctoral Fellowship and through the support of the ID# 62312 grant from the John Templeton Foundation, as part of the ‘The Quantum Information Structure of Spacetime’ Project (QISS).